The following information has been taken from the books of Lobelia Traders, a general trading shop in Braamfontein, for the month of April 2020: Lobelia Traders is a registered VAT vendor and trades only with registered VAT Vendors. The business charges 15% VAT on all its sales. TRANSACTION JOURNAL AMOUNT EXCLUDING VAT VAT AMOUNT AMOUNT INCLUDING VAT Cash sales CBR 35 000 A  B Advertising paid CBP C D 920 Goods purchased for cash CBP E 750 F Drawings of Petty Cash GJ 500 G H Credit sales DJ I 1 200 J    Q.3.2.1  Calculate the amounts for A to J (10) Q.3.2.2  Calculate the Input VAT Total (2½) Q.3.2.3  Calculate  the Output VAT total (2½) Q.3.2.4 Calculate the Vat payable or refundable. Clearly indicate if it is payable or refundable and if it is an asset or liability  (5)


1
Expert's answer
2021-05-04T07:24:07-0400

Answers:

(i) Determination of missing values:

Cash sales;

A (VAT amount); 15% of 35000 = 5250

B (Amount including VAT); 35000 + 5250 = 40250


Advertising paid:

C (Amount excluding VAT); (100% * 920)/115 = 800

D (VAT amount) = 920 - 800 or 15% of 800 = 120


Goods purchased for cash:

E (Amount excluding VAT); (100% * 750)/15% = 5000

F (Amount including VAT); 5000 + 750 = 5750


Drawings of petty cash:

G (VAT amount); 15% of 500 = 75

H (Amount including VAT); 500 + 75 = 575


Credit Sales:

I (Amount excluding VAT); (100% * 1200)/15% = 8000

J (Amount including VAT); 8000 + 1200 = 9200


(ii) Determination of VAT Payable or Refundable:

VAT Payable = Total Output VAT - Total Input VAT

Total output VAT; 5250 + 120 0 = 6450

Total input VAT; 120 + 750 + 75 = 945


"\therefore" VAT Payable ; 6450 - 945 = 5,505 (Liability)